OpenStudy (anonymous):

What is the solution of: √2x + 4 - √x=2 A. x=0 B. x=16 and x=-16 C. x=0 and c = 16 D. x=0 and x=-16

OpenStudy (anonymous):

well it's obivously not 0, so....

OpenStudy (anonymous):

Ok so i can cancel A out

OpenStudy (anonymous):

well since all answers except b contain 0 as an answer, you can prettymuch cancel everyting except b out

OpenStudy (anonymous):

Oh ok thank you

OpenStudy (anonymous):

yes, i'm sorry it's so easy, but it's quite obvious that when you cancel out all of the x by making them 0, you'll only have 4=2 left for your equation, which is obviously not true.

OpenStudy (anonymous):

\[\sqrt{2x}+4-\sqrt{x}=2\] ...if this is the problem.. b is not correct.. none of the answers are.

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

actually nameless is right, there are no solutions, whoever gave you this question made a mistake.

OpenStudy (anonymous):

It was a question from my test

OpenStudy (anonymous):

Did the test ask \( \sqrt{2x}+4-\sqrt{x}=2 \) or \( \sqrt{2}x+4-\sqrt{x}=2 \) and was it a test on complex numbers?

OpenStudy (anonymous):

because the first equation is impossable to solve, and the second equation is only possible if you use complex numbers, none of the above answers is even close to correct.

OpenStudy (anonymous):

so whoever made that test made a mistake. And you should tell him/her.

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

\[\large \sqrt{2*16}+4-\sqrt{16}=4\sqrt{2}\]

OpenStudy (anonymous):

not the same as 2